# Adversarial Poisoning Attack on Quantum Machine Learning Models

### v2

@@ I Introduction
Quantum computing is rapidly progressing, with companies like Atom Computing and IBM recently unveiling the largest quantum processors ever developed, boasting 1,225 and 1,121 qubits, respectively [1, 2]. The significant interest in quantum computing stems from its potential to offer substantial computational speedups over classical computers for certain problems. Researchers have already begun leveraging current noisy intermediate-scale quantum (NISQ) machines to demonstrate practical utility in this pre-fault-tolerant era [3]. Within this emerging field, quantum machine learning (QML) has also gained considerable attention, merging the power of quantum computing with classical machine learning algorithms. QML explores the potential to improve learning algorithms by leveraging unique quantum properties like superposition and entanglement, opening new horizons in computational speed and capability. Several QML models have been explored, with quantum neural networks (QNNs) [4, 5, 6] standing out as the most popular model, mirroring the structure and function of classical neural networks within a quantum framework.
@@
However, the rapidly evolving field of quantum computing brings not only unique computational capabilities but also a variety of new security challenges. Several recent works have explored the security aspects of quantum circuits in current NISQ-era. Authors in [7] introduced a power side-channel attack on quantum computer controllers, demonstrating how power trace analysis could be used to reconstruct quantum circuits. Researchers have also begun analyzing the effectiveness of adversarial attacks on QML models, such as model stealing attacks [8], where an adversary can replicate the functionality of a black-box QNN using only input-output queries. Novel defenses have been proposed to mitigate such attack scenarios [9]. Additionally, recent works have highlighted the threat of backdoor attacks [10, 11] on QNNs, revealing them to be a significant concern for QML models. Another critical threat to QML models in current cloud settings, considered the greatest concern for classical ML deployment [12, 13], is data poisoning attacks, which have not yet been explored in the context of QML.
@@
Data poisoning attacks, first proposed in [14], are a class of adversarial attacks aimed at manipulating a victim model by either injecting poisoned samples into the dataset or perturbing a subset of the training data $\mathcal{D}_{tr}$. A detailed description of data poisoning attacks and their sub-classes can be found in Section II-B. These attacks have been extensively studied in the classical domain, where most approaches rely on a deep understanding of the victim’s training procedure (unrealistic) and gradient optimization techniques to generate minimally perturbed poisoned samples (costly). Furthermore, such methods often lack robustness to the noise present in modern NISQ devices, highlighting the need for a fresh perspective on data poisoning attacks in QML. Building on how data poisoning initially focused predominantly on label-flipping attacks, we take a first step in exploring indiscriminate data poisoning in QML by proposing a novel label-flipping technique based on intra-class state similarity in quantum Hilbert space.
@@
In this work, we propose a quantum indiscriminate data poisoning attack, QUID, using a novel label-flipping strategy. QUID assigns poisoned labels to a subset of the training data such that the encoded state is furthest from its true class. Unlike other classical label-flipping [15, 16] that require separately training the victim model to find adversarial labels, QUID avoids this costly training process while being noise-robust, as it can effectively determine poisoned labels by executing the encoding circuit on noisy hardware. Our main contributions in this work are:
@@
  • Evaluated the performance of QUID against random flipping and randomly poisoned samples across various QNN architectures and datasets in both noiseless and noisy environments, along with testing against a state-of-the-art defense technique, demonstrating its effectiveness.
@@
II Background
II-A Quantum Neural Network (QNN)
A QNN comprises three key components: (i) a classical-to-quantum data encoding circuit, (ii) a parameterized quantum circuit (PQC) optimized for specific tasks, and (iii) measurement operations. For continuous variables, angle encoding is widely used, mapping classical features to qubit rotations along a chosen axis [5]. Since qubit states repeat every $2\pi$, features are typically scaled to 00 to $2\pi$ (or $-\pi$ to $\pi$) during preprocessing. In this study, we use $RZ$ gates for encoding classical features into quantum states.
Figure 1: Architecture of a 4-qubit hybrid QNN. Classical features are encoded as angles of quantum rotation gates ($R_{Z}$). The PQC transforms the encoded states to explore the search space and entangle features. Measured expectation values are then fed into a classical linear layer for the final prediction.
@@ II-A Quantum Neural Network (QNN)
II-B Data Poisoning Attack
Data poisoning attacks [14, 17] are a class of adversarial attacks aimed at manipulating data by either tampering the current samples/labels or injecting poisoned samples to compromise the integrity and/or availability of the model. Data poisoning attacks can be broadly classified into three main categories: i) Targeted attacks, which preserve the functionality and behavior of the system for general users while causing misclassification for specific target samples; ii) Indiscriminate attacks, which aim to degrade model performance on test data; and iii) Backdoor attacks, which, for any test sample containing a specific pattern, known as the backdoor trigger, induce misclassification without affecting the classification of clean test samples. In this work, we focus on indiscriminate data poisoning attacks on QNNs which can further be divided into; a) Label-flip poisoning, where adversary does not perturb the features but only flips label of a subset of the dataset, b) Bilevel poisoning, where attacker can obfuscate both the training samples and their labels and c) Clean-label poisoning where attackers only slightly perturb each sample to preserve it’s class i.e. performing a clean-label attack. Here, we propose a quantum state proximity based label-poisoning attack.
TABLE I: Comparison with Related Classical Attacks.
@@ II-B Data Poisoning Attack
[16]
[18]
[19]
QUID (ours)
II-C Related Works and Limitations
Although there has been work on backdoor attacks in QNNs [11, 10], some of which modify QNN architecture to perform indiscriminate attacks, no research has yet explored the viability of data poisoning attacks on QNNs. Current state-of-the-art (SOTA) classical indiscriminate data poisoning attacks [19, 17] are gradient-based and assume that the adversary has access to the training data ($\mathcal{D}_{tr}$), the target model (white-box access), the training procedure (loss function, optimization process, hyperparameters, etc.), and sometimes even the test data [18]. QUID, on the other hand, only requires access to $\mathcal{D}_{tr}$ and gray-box access to the QNN circuit, i.e., only the encoding circuit of the QNN. This is a more realistic and efficient framework considering the high cost associated with trainings QNN on NISQ devices.
Apart from relying on extensive knowledge and expensive optimization procedures to poison the dataset, most classical SOTA attacks ensure that the poisoned samples, i.e., $\{x_{i}\mid(x_{i},y_{i})\in\mathcal{D}_{p}\subset\mathcal{D}_{tr}\}$, are minimally perturbed to avoid detection. For example, the Gradient Canceling (GC) attack [19] aims to manipulate the model’s training process by injecting slightly perturbed poisoned data into $\mathcal{D}_{tr}$, with the goal of disrupting learning by “canceling out” the gradients that guide parameter updates during training. However, due to the noisy nature of current quantum hardware, these small perturbations, effective in classical deep learning, are masked by the inherent quantum device noise, rendering them ineffective (Table I). Studies have even shown that inherent noise makes QNNs adversarially robust [20, 21]. Consequently, there is a need to reevaluate data poisoning attacks in the context of QML.
III Adversarial Framework
III-A Attack Model
@@
Figure 2: Overview of QUID’s label-poisoning technique. The adversary extracts the encoding circuit from the QNN and uses it to compute the corresponding output density matrix ($\sigma$) for a portion ($\varepsilon$) of the training data $\mathcal{D}_{train}$. It then calculates the matrix distance between $\sigma$ and the remaining samples in the dataset ($\rho_{i}$), assigning the class with the maximum distance ($\max d_{\sigma\rho}$).
III-B Adversary Objective
@@ IV Proposed Methodology
Figure 3: Figure illustrating the accuracy of labels generated based on $\min$ intra-class state distances compared to the true labels for the reduced MNIST-10 dataset with a latent dimension of $d=8$, evaluated under both noiseless and noisy environments.
IV-A Intra-Class Encoder State Similarity (ESS)
@@
TABLE II: Performance comparison of label assignment based on $\min$ intra-class distance with respect to true labels, across various 4-class datasets and distance metrics. Frobenius norm appears to be the most optimal choice due to its superior performance and efficient computational cost.
@@
Fig. 3 illustrates how accurately we can identify classes in a portion (50%) of a reduced MNIST-10 dataset using intra-class state similarity. Under both noiseless and noisy scenarios, the Frobenius norm outperforms the HS and Trace norms in determining the correct classes. Table II further presents the performance of different 4-class datasets under noisy conditions. Here too, we observe that, on average, using the Frobenius norm to calculate the distance between density matrices results in better performance while requiring considerably less computation time, making it the optimal choice of distance metric. Furthermore, we believe ESS can be used to estimate the performance of encoding methods and to estimate the separability of datasets when mapped to $\mathcal{H}$ without the need for training the entire QNN from scratch.
IV-B QUID Poisoning Strategy
@@
where $\theta$ represents the set of parameters of $\mathcal{F}$. We use the intra-class quantum state similarity concept to find adversarial labels which leads to high performance degradation. Given that for any dataset, intra-class distances and inter-class distances in $\mathcal{H}$ should be low and high respectively, we modify the labels of a subset ($\varepsilon$) of the training data $\mathcal{D}_{tr}$ which leads to high intra-class distances. Specifically, given the training dataset $\mathcal{D}_{tr}$, we first split it into clean $\mathcal{D}_{c}$ and poisoned dataset $\mathcal{D}_{p}$ based on the poison ratio $\varepsilon$. For samples $\{x_{i}\mid(x_{i},y_{i})\in\mathcal{D}_{p}\}$, find density matrix of the encoded states $\phi(x_{i}):x_{i}\in\mathcal{D}_{p}$, calculate distance between the encoded states $\phi(x_{i}):x_{i}\in\mathcal{D}_{c}$ and assign the class label with $\mathbf{\max}$ distance. An algorithmic description of QUID’s label poisoning technique is described in Algorithm 1. Essentially, we propose a greedy algorithm to determine optimal labels, in contrast to [16] which relies on a brute force training method involving iterating through all possible labels for each sample and selecting the label that results in the maximum loss.
@@
V Evaluation
V-A Setup
Device: We conducted all noiseless experiments on Pennylane’s [22] lightning.qubit device using the “Adjoint” differentiation method to calculate gradients. For noisy training, we employed the default.mixed device and introduced Amplitude Damping and Depolarizing Channel noise after each gate in the circuit, with a high error probability of $p=0.05$ (unless stated otherwise) to effectively simulate the noisy nature of current hardware. Considering that generic methods for calculating gradients, such as backpropagation, are not viable on actual quantum devices, we used the Simultaneous Perturbation Stochastic Approximation (SPSA) method for gradient calculation in our QNN training. This approach generates noisy gradients, providing a more realistic representation of the performance expected from real quantum hardware and ensuring that our experimental results align more closely with practical quantum computing environments.
Training: For evaluation, we used PQC-1, 6 and 8 to build our QNNs (PQC-$x$ represents circuit-$x$ in [23]), initialized with random weights. In most of the experiments we used a 4-qubit or 8-qubit QNN and angle encoding to encode classical features, consistent with recent QML works [24, 25]. For training the QNNs we used CrossEntropyLoss() as our classification loss function. The hyperparameters used for training are; Epochs: 30, learning_rate ($\eta$): 0.01, batch_size = 32 and optimizer: Adam. All training are done on an Intel Core-i9 13900K CPU with 64GB of RAM.
Dataset: Since NISQ devices struggle with large images due to limited qubits, high error rates and complex data encoding, similar to recent works [26, 25], we conduct all experiments using a reduced feature set of MNIST, Fashion, Kuzushiji and Letters datasets with latent dimension $d=8$ (from original 28$\times$28 image) generated using a convolutional autoencoder [27]. Thus, for each dataset, we create a smaller 4-class dataset from these reduced feature sets i.e., MNIST-4 (class 0, 1, 2, 3), Fashion-4 (class 6, 7, 8, 9), Kuzushiji-4 (class 3, 5, 6, 9) and Letters-4 (class 1, 2, 3, 4) with each having 1000 samples (700 for training and 300 for testing). Since each of these datasets is of dimension $d=8$, we encode 2 features per qubit. Number of shots/trials is set to 1000 for all experiments.
Figure 4: Performance of angle encoding (4 features/qubit) and amplitude encoding when using ESS to determine embedding circuit performance under varying noise levels.
TABLE III: Test accuracies for label-flip (LF) and bi-level poisoning attacks on different datasets, where poisoned samples (PS) $x_{i}\in\mathcal{D}_{p}$ are randomly generated. QUID achieves maximal performance degradation when only the labels are poisoned.

### v3

# Adversarial Data Poisoning Attacks on Quantum Machine Learning in the NISQ Era

@@ 1. Introduction
Quantum computing is rapidly progressing, with companies like Atom Computing and IBM recently unveiling the largest quantum processors ever developed, boasting 1,225 and 1,121 qubits, respectively (Computing, 2023; Gambetta, 2023). The significant interest in quantum computing stems from its potential to offer substantial computational speedups over classical computers for certain problems. Researchers have already begun leveraging current noisy intermediate-scale quantum (NISQ) machines to demonstrate practical utility in this pre-fault-tolerant era (Kim et al., 2023). Within this emerging field, quantum machine learning (QML) has also gained considerable attention, merging the power of quantum computing with classical machine learning algorithms. QML explores the potential to improve learning algorithms by leveraging unique quantum properties like superposition and entanglement, opening new horizons in computational speed and capability. Several QML models have been explored, with quantum neural networks (QNNs) (Schuld et al., 2014; Abbas et al., 2021; Schuld et al., 2020) standing out as the most popular model, mirroring the structure and function of classical neural networks within a quantum framework.
@@
However, the rapidly evolving field of quantum computing brings not only unique computational capabilities but also a variety of new security challenges. Several recent works have explored the security aspects of quantum circuits in current NISQ-era. Authors in (Erata et al., 2024) introduced a power side-channel attack on quantum computer controllers, demonstrating how power trace analysis could be used to reconstruct quantum circuits. Researchers have also begun analyzing the effectiveness of adversarial attacks on QML models, such as model stealing attacks (Kundu et al., 2024), where an adversary can replicate the functionality of a black-box QNN using only input-output queries. Novel defenses have been proposed to mitigate such attack scenarios (Kundu and Ghosh, 2024). Additionally, recent works have highlighted the threat of backdoor attacks (Chu et al., 2023a, b) on QNNs, revealing them to be a significant concern for QML models. Another critical threat to QML models in current cloud settings, considered the greatest concern for classical ML deployment (Grosse et al., 2023; Kumar et al., 2020), is data poisoning attacks, which have not yet been explored in the context of QML.
@@
Data poisoning attacks, first proposed in (Biggio et al., 2012), are a class of adversarial attacks aimed at manipulating a victim model by either injecting poisoned samples into the dataset or perturbing a subset of the training data $\mathcal{D}_{tr}$. A detailed description of data poisoning attacks and their sub-classes can be found in Section 2.2. These attacks have been extensively studied in the classical domain, where most approaches rely on a deep understanding of the victim’s training procedure (unrealistic) and gradient optimization techniques to generate minimally perturbed poisoned samples (costly). Furthermore, such methods often lack robustness to the noise present in modern NISQ devices, highlighting the need for a fresh perspective on data poisoning attacks in QML. Building on how data poisoning initially focused predominantly on label-flipping attacks, we take a first step in exploring indiscriminate data poisoning in QML by proposing a novel label-flipping technique based on intra-class state similarity in quantum Hilbert space.
@@
In this work, we propose a quantum indiscriminate data poisoning attack, QUID, using a novel label-flipping strategy. QUID assigns poisoned labels to a subset of the training data such that the encoded state is furthest from its true class. Unlike other classical label-flipping (Taheri et al., 2020; Paudice et al., 2019) that require separately training the victim model to find adversarial labels, QUID avoids this costly training process while being noise-robust, as it can effectively determine poisoned labels by executing the encoding circuit on noisy hardware. Our main contributions in this work are:
@@
  • Evaluated the performance of QUID against random flipping and randomly poisoned samples across various QNN architectures and datasets in both noiseless and noisy environments, along with testing against a state-of-the-art defense technique (Levine and Feizi, 2021), demonstrating its effectiveness.
@@
2. Background
2.1. Quantum Neural Network (QNN)
A QNN comprises three key components: (i) a classical-to-quantum data encoding circuit, (ii) a parameterized quantum circuit (PQC) optimized for specific tasks, and (iii) measurement operations. For continuous variables, angle encoding is widely used, mapping classical features to qubit rotations along a chosen axis (Abbas et al., 2021). Since qubit states repeat every $2\pi$, features are typically scaled to 00 to $2\pi$ (or $-\pi$ to $\pi$) during preprocessing. In this study, we use $RZ$ gates for encoding classical features into quantum states.
Figure 1. Architecture of a 4-qubit hybrid QNN. Classical features are encoded as angles of rotation gates ($R_{Z}$). The PQC transforms the encoded states to explore the search space and entangle features. Measured expectation values are then fed into a classical linear layer for the final prediction.
@@ 2.1. Quantum Neural Network (QNN)
2.2. Data Poisoning Attack
Data poisoning attacks (Biggio et al., 2012; Lu et al., 2022) are a class of adversarial attacks aimed at manipulating data by either tampering the current samples/labels or injecting poisoned samples to compromise the integrity and/or availability of the model. Data poisoning attacks can be broadly classified into three main categories: i) Targeted attacks, which preserve the functionality and behavior of the system for general users while causing misclassification for specific target samples; ii) Indiscriminate attacks, which aim to degrade model performance on test data; and iii) Backdoor attacks, which, for any test sample containing a specific pattern, known as the backdoor trigger, induce misclassification without affecting the classification of clean test samples. In this work, we focus on indiscriminate data poisoning attacks on QNNs which can further be divided into; a) Label-flip poisoning, where adversary does not perturb the features but only flips label of a subset of the dataset, b) Bilevel poisoning, where attacker can obfuscate both the training samples and their labels and c) Clean-label poisoning where attackers only slightly perturb each sample to preserve it’s class i.e. performing a clean-label attack. Here, we propose a quantum state proximity based label-poisoning attack.
Table 1. Comparison with related classical attacks.
@@ 2.2. Data Poisoning Attack
(Paudice et al., 2019)
(Koh et al., 2022)
(Lu et al., 2023)
Ours
2.3. Related Works and Limitations
Although there has been work on backdoor attacks in QNNs (Chu et al., 2023b, a), some of which modify QNN architecture to perform indiscriminate attacks, no research has yet explored the viability of data poisoning attacks on QNNs. Current state-of-the-art (SOTA) classical indiscriminate data poisoning attacks (Lu et al., 2023, 2022) are gradient-based and assume that the adversary has access to the training data ($\mathcal{D}_{tr}$), the target model (white-box access), the training procedure (loss function, optimization process, hyperparameters, etc.), and sometimes even the test data (Koh et al., 2022). QUID, on the other hand, only requires access to $\mathcal{D}_{tr}$ and gray-box access to the QNN circuit, i.e., only the encoding circuit of the QNN. This is a more realistic and efficient framework considering the high cost associated with trainings QNN on NISQ devices.
Apart from relying on extensive knowledge and expensive optimization procedures to poison the dataset, most classical SOTA attacks ensure that the poisoned samples, i.e., $\{x_{i}\mid(x_{i},y_{i})\in\mathcal{D}_{p}\subset\mathcal{D}_{tr}\}$, are minimally perturbed to avoid detection. For example, the Gradient Canceling (GC) attack (Lu et al., 2023) aims to manipulate the model’s training process by injecting slightly perturbed poisoned data into $\mathcal{D}_{tr}$, with the goal of disrupting learning by “canceling out” the gradients that guide parameter updates during training. However, due to the noisy nature of current quantum hardware, these small perturbations, effective in classical deep learning, are masked by the inherent quantum device noise, rendering them ineffective (Table 1). Studies have even shown that inherent noise makes QNNs adversarially robust (Du et al., 2021; Wang et al., 2021). Consequently, there is a need to reevaluate data poisoning attacks in the context of QML.
While label-flipping attacks exist in classical ML, QUID’s novelty lies in adapting this idea to QML setting by leveraging quantum properties, such as state similarity in Hilbert space. To date, no prior work has focused on poisoning attacks or defenses specific to QML models, which is also why quantum-specific defenses remain unexplored. We acknowledge this limitation in Section 5.2.
3. Adversarial Framework
3.1. Attack Model
@@
Figure 2. Overview of QUID’s label-poisoning technique. The adversary extracts the encoding circuit from the QNN and uses it to compute the corresponding output density matrix ($\sigma$) for a portion ($\varepsilon$) of the training data $\mathcal{D}_{train}$. It then calculates the matrix distance between $\sigma$ and the remaining samples in the dataset ($\rho_{i}$), assigning the class with the maximum distance ($\max d_{\sigma\rho}$).
3.2. Adversary Objective
@@ 4. Proposed Methodology
Figure 3. Figure illustrating the accuracy of labels generated based on $\min$ intra-class state distances compared to the true labels for the reduced MNIST-10 dataset with a latent dimension of $d=8$, evaluated under both noiseless and noisy environments.
4.1. Intra-Class Encoder State Similarity (ESS)
@@
Table 2. Performance comparison of label assignment based on $\min$ intra-class distance with respect to true labels, across various 4-class datasets and distance metrics. Frobenius norm appears to be the most optimal choice due to its superior performance and efficient computational cost.
@@
Fig. 3 illustrates how accurately we can identify classes in a portion (50%) of a reduced MNIST-10 dataset using intra-class state similarity. Under both noiseless and noisy scenarios, the Frobenius norm outperforms the HS and Trace norms in determining the correct classes. Table 2 further presents the performance of different 4-class datasets under noisy conditions. Here too, we observe that, on average, using the Frobenius norm to calculate the distance between density matrices results in better performance while requiring considerably less computation time, making it the optimal choice of distance metric. Furthermore, we believe ESS can be used to estimate the performance of encoding methods and to estimate the separability of datasets when mapped to $\mathcal{H}$ without the need for training the entire QNN from scratch.
4.2. QUID Poisoning Strategy
@@
where $\theta$ represents the set of parameters of $\mathcal{F}$. We use the intra-class quantum state similarity concept to find adversarial labels which leads to high performance degradation. Given that for any dataset, intra-class distances and inter-class distances in $\mathcal{H}$ should be low and high respectively, we modify the labels of a subset ($\varepsilon$) of the training data $\mathcal{D}_{tr}$ which leads to high intra-class distances. Specifically, given the training dataset $\mathcal{D}_{tr}$, we first split it into clean $\mathcal{D}_{c}$ and poisoned dataset $\mathcal{D}_{p}$ based on the poison ratio $\varepsilon$. For samples $\{x_{i}\mid(x_{i},y_{i})\in\mathcal{D}_{p}\}$, find density matrix of the encoded states $\phi(x_{i}):x_{i}\in\mathcal{D}_{p}$, calculate distance between the encoded states $\phi(x_{i}):x_{i}\in\mathcal{D}_{c}$ and assign the class label with $\mathbf{\max}$ distance. An algorithmic description of QUID’s label poisoning technique is described in Algorithm 1. Essentially, we propose a greedy algorithm to determine optimal labels, in contrast to (Paudice et al., 2019) which relies on a brute force training method involving iterating through all possible labels for each sample and selecting the label that results in the maximum loss.
@@
5. Evaluation
5.1. Experimental Setup
Device: We conducted all noiseless experiments on Pennylane’s (Bergholm et al., 2018) lightning.qubit device using the “Adjoint” differentiation method to calculate gradients. For noisy training, we employed the default.mixed device and introduced Amplitude Damping and Depolarizing Channel noise after each gate in the circuit, with a high error probability of $p=0.05$ (unless stated otherwise) to effectively simulate the noisy nature of current hardware. Considering that generic methods for calculating gradients, such as backpropagation, are not viable on actual quantum devices, we used the Simultaneous Perturbation Stochastic Approximation (SPSA) method for gradient calculation in our QNN training. This approach generates noisy gradients, providing a more realistic representation of the performance expected from real quantum hardware and ensuring that our experimental results align more closely with practical quantum computing environments.
Training: For evaluation, we used PQC-1, 6 and 8 to build our QNNs (PQC-$x$ represents circuit-$x$ in (Sim et al., 2019)), initialized with random weights. In most of the experiments we used a 4-qubit or 8-qubit QNN and angle encoding to encode classical features, consistent with recent QML works (Wang et al., 2022a; Anagolum et al., 2024). For training the QNNs we used CrossEntropyLoss() as our classification loss function. The hyperparameters used for training are; Epochs: 30, learning_rate ($\eta$): 0.01, batch_size = 32 and optimizer: Adam. All training are done on an Intel Core i9-13900K CPU with 64GB of RAM.
Dataset: Since NISQ devices struggle with large images due to limited qubits, high error rates and complex data encoding, similar to recent works (Wang et al., 2022b; Anagolum et al., 2024), we conduct all experiments using a reduced feature set of MNIST, Fashion, Kuzushiji and Letters datasets with latent dimension $d=8$ (from original 28$\times$28 image) generated using a convolutional autoencoder (Alam et al., 2021). Thus, for each dataset, we create a smaller 4-class dataset from these reduced feature sets i.e., MNIST-4 (class 0, 1, 2, 3), Fashion-4 (class 6, 7, 8, 9), Kuzushiji-4 (class 3, 5, 6, 9) and Letters-4 (class 1, 2, 3, 4) with each having 1000 samples (700 for training and 300 for testing). Since each of these datasets is of dimension $d=8$, we encode 2 features per qubit. Number of shots/trials is set to 1000 for all experiments.
Figure 4. Performance of angle encoding (4 features/qubit) and amplitude encoding when using ESS to determine embedding circuit performance under varying noise levels.
Table 3. Test accuracies for label-flip (LF) and bi-level poisoning attacks on different datasets, where poisoned samples (PS) $x_{i}\in\mathcal{D}_{p}$ are randomly generated. QUID achieves maximal performance degradation when only the labels are poisoned.